# Coarse obstructions to cocompact cubulation

### v1

@@ 1. Introduction
A tubular group admits a coarse median if and only if it is virtually compact special.
In [Wis14b], Wise characterised which tubular groups admit free actions on CAT(0) cube complexes. These cubulations were further investigated by Woodhouse [Woo16, Woo18]. Wise also showed that all cocompactly cubulated tubular groups are virtually special, and we use this in the proof of Theorem E.
@@
We are indebted to Naomi Andrew for several very useful discussions about free-by-cyclic groups. We are very grateful to Mark Hagen and Monika Kudlinska for helpful conversations about train tracks. We thank Jean Pierre Mutanguha for comments on an earlier version. We also thank the organisers of the 2023 thematic program on geometric group theory in Montreal, where this work began.
@@ 2.3. Coarse medians
Whilst the fact that $\hat{X}$ is a median metric space of finite rank is already useful, the exact control on ranks will be refined in Proposition 3.2 below, which shows that if $\hat{X}$ is an asymptotic cone of a coarse median space $X$, then $\operatorname{rk}X=\operatorname{rk}\hat{X}$. This justifies the similarity in notation between the rank of a coarse median space $X$ and that of a median algebra.
@@ 3. Dimension bounds
First we establish a technical lemma that allows one to pass from having bilipschitz cubes in the asymptotic cone of a coarse median space $X$ to having large uniformly quasiisometrically embedded cubes in $X$. This is similar to part of the argument of Lemma 2.9.
3.1 Lemma.
Let $X$ be a coarse median space. For each $d$ there is a constant $q$ such that the following holds. If there is a median monomorphism of $\{0,1\}^{d}$ in an asymptotic cone $\hat{X}$ of $X$, then there are $q$–quasimedian $q$–quasiisometries $[0,n]^{d}\to X$ for all $n$.
Proof.
Suppose that $f:\{0,1\}^{d}\to\hat{X}=\lim_{\omega}(X,\lambda_{n}\operatorname{\mathsf{d}})$ is a median monomorphism. After rescaling $\hat{X}$, we can assume that points in the image of $f$ lie at distance at least 1 from each other. For each $p\in\{0,1\}^{d}$, let $(x_{n}^{p})$ be a sequence in $X$ such that $(x_{n}^{p})=f(p)$. Let $h=h_{2^{d}}$ be given by the definition of a coarse median space, and let $t=t(d)$, $r=r(d)$ be given by [Bow19, Lem. 7.2]. We are free to assume (for convenience) that $h,t,r\geq 1$. Let $\varepsilon>0$ be such that $\frac{1-\varepsilon}{\varepsilon}>10htr$. We have
$N_{\varepsilon}\,=\,\big{\{}m\in\mathbf{N}\,:\,\{x_{m}^{p}:p\in\{0,1\}^{d}\}\text{ is a }\frac{1-\varepsilon}{\lambda_{m}}\text{--separated }\frac{\varepsilon}{\lambda_{m}}\text{--quasisubalgebra}\big{\}}\,\in\,\omega.$
In particular, $N_{\varepsilon}$ contains arbitrarily large elements.
Given $m\in N_{\varepsilon}$, if we let $\kappa_{m}=\frac{\varepsilon}{\lambda_{m}}$, then we have that $\{x^{p}_{m}\}$ is a $10\kappa_{m}htr$–separated $\kappa_{m}$–quasisubalgebra consisting of $2^{d}$ points. According to [Bow19, Lem. 7.2, 9.1], there is a median algebra $\Pi_{m}$ of cardinality at most $2^{2^{d}}$ such that the following hold.
  • $\Pi_{m}$ is a $d$–dimensional $h$–quasisubalgebra.
  • There is a median epimorphism $\pi_{m}:\Pi_{m}\to\{0,1\}^{d}$ such that $\operatorname{\mathsf{d}}(q,x^{\pi_{m}(q)}_{m})\leq\kappa_{m}r$.
  • $\Pi_{m}$ is “$s$–straight”, where $s$ depends only on the coarse median parameters.
By the second of these properties, each $\Pi_{m}$ consists of a union of clusters $\{\pi_{m}^{-1}(x_{m}^{p})\}$, each of which is at a distance of at least $5\kappa_{m}htr$ from all others. We deduce that $\Pi_{m}$ contains a $5\kappa_{m}htr$–separated $h$–quasisubalgebra $Q_{m}$ isomorphic to a $d$–cube.
By the first and third items above, each $Q_{m}$ meets the assumptions of [Bow19, Lem. 9.3]. For each $m$, an application of that lemma now provides a uniform quasimedian quasiisometry from a product of $d$ intervals of length at least $5\kappa_{m}htr$ to the coarse-median hull of $Q_{m}$. The lemma follows by restricting these maps to subcubes. ∎
Using this, we can obtain the following characterisation of the infimal rank of a coarse median space. By the quasiflat rank of a metric space $X$, we mean the supremal integer $\operatorname{qf.rk}X$ for which there is a quasiisometric embedding $\mathbf{R}^{\operatorname{qf.rk}X}\to X$.
@@
Let $X$ be a coarse median space, and let $\hat{X}$ be an asymptotic cone of $X$, which is a median algebra of rank $\operatorname{rk}\hat{X}$. The following quantities agree.
@@
  • The supremal $d$ such that there is a bilipschitz, median embedding of $[0,1]^{d}$ in $\hat{X}$.
@@
If $X$ is proper and has cocompact isometry group, then we also have $\operatorname{rk}X=\operatorname{qf.rk}X$. In this case, there exists a quasimedian quasiisometric embedding $\mathbf{R}^{\operatorname{rk}X}\to X$, for some coarse median on $X$ realising $\operatorname{rk}X$.
@@
The agreement of $\operatorname{sepdim}\hat{X}$ with $\operatorname{rk}\hat{X}$ is given by [Hae16, Cor. 3.7]. If $\operatorname{rk}\hat{X}\geq d$, then by [Hae16, Thm G] there is a bilipschitz, median embedding of $[0,1]^{d}$ in $\hat{X}$. By Lemma 3.1, this implies the existence of arbitrarily large uniform quasicubes in $X$, which shows that $\operatorname{rk}X\geq d$. Thus $\operatorname{rk}X\geq\operatorname{rk}\hat{X}$. The converse is [Bow13, Thm 2.3].
If $\operatorname{qf.rk}X\geq d$, then $\hat{X}$ contains a bilipschitz copy of $\mathbf{R}^{d}$ and hence $\operatorname{sepdim}\hat{X}\geq d$. By the above, we therefore have $\operatorname{rk}X\geq\operatorname{qf.rk}X$. It remains to show that if $X$ is proper and has cocompact isometry group, then $X$ admits a quasimedian quasiisometric embedding of $\mathbf{R}^{\operatorname{rk}X}$, for this in particular implies that $\operatorname{qf.rk}X\geq\operatorname{rk}X$.
Suppose that $X$ is proper and has cocompact isometry group. Fix a basepoint $x_{0}\in X$. Let $\mu$ be a coarse median on $X$ with rank equal to $\operatorname{rk}X$. For $n\in\mathbf{N}$, let $Q_{n}$ be a uniform quasicube in $X$ of dimension $d=\operatorname{rk}X$ and with diameter at least $n$. Write $z_{n}$ for the central point of $Q_{n}$. By cocompactness, we can translate $z_{n}$ into a fixed compact set $C$ containing $x_{0}$ by an isometry $g_{n}$. This gives uniform quasicubes in the sequence of coarse median spaces $(X,g_{n}\mu)$ that are all centred in the compact set $C$.
Now take an (unrescaled) ultralimit. Since $X$ is proper, we have $X=\lim_{\omega}B(x_{0},m)$, where $B(x_{0},m)$ is the ball of radius $m$ centred on $x_{0}$. Choosing $m_{n}$ so that $g_{n}Q_{n}\subset B(x_{0},m_{n})$, we get that $Q=\lim_{\omega}g_{n}Q_{n}\subset X$, because $g_{n}Q_{n}$ is centred in the fixed compact set $C$. Because the $Q_{n}$ are uniform quasiisometric embeddings of increasingly large $d$–cubes, $Q$ is a quasiisometric embedding of $\mathbf{R}^{d}$.
@@ 6.2. Coarse median tubular groups
Let $G$ be a rank-$2$ tubular group. If $G$ admits a coarse median, then $G$ is virtually compact special.
@@
According to [Bow13, Cor. 8.3], if $G$ admits a coarse median then it has a quadratic isoperimetric function, and so Theorem 6.8 shows that $G$ must be undistorted. Moreover, $G$ has geometric dimension two since it is a graph of two dimensional groups (see [SW79, Prop. 3.6]), so Theorem 3.6 shows that $\operatorname{rk}G\leq 2$. In particular, Theorem 4.3 implies that $G$ cannot have a quasiisometrically embedded 2–RBF. By Lemma 6.9, this means that no vertex group of $G$ can have more than two commensurability classes of incident edge groups. Since $G$ is undistorted, it contains no Baumslag–Solitar subgroups $\operatorname{BS}(m,n)$ with $m\neq\pm n$. Consequently, [Wis14b, Cor 5.10, 5.9] implies that $G$ is virtually compact special. ∎

### v2

@@ 1. Introduction
A tubular group admits a coarse median if and only if it is cocompactly cubulated and virtually compact special.
In [Wis14b], Wise characterised which tubular groups admit free actions on CAT(0) cube complexes. These cubulations were further investigated by Woodhouse [Woo16, Woo18]. Wise also showed that all cocompactly cubulated tubular groups are virtually special.
@@
We are indebted to Naomi Andrew for several very useful discussions about free-by-cyclic groups. We are very grateful to Mark Hagen and Monika Kudlinska for helpful conversations about train tracks. We thank Jason Behrstock and Jean Pierre Mutanguha for comments on an earlier version. We also thank the organisers of the 2023 thematic program on geometric group theory in Montreal, where this work began.
@@ 2.3. Coarse medians
Whilst the fact that $\hat{X}$ is a median metric space of finite rank is already useful, the exact control on ranks will be refined in Proposition 3.1 below, which shows that if $\hat{X}$ is an asymptotic cone of a proper cocompact coarse median space $X$, then $\operatorname{rk}X=\operatorname{rk}\hat{X}$. This justifies the similarity in notation between the rank of a coarse median space $X$ and that of a median algebra.
@@ 3. Dimension bounds
The goal of this section is to upper bound the minimal rank of a coarse median space in terms of its geometry. By the quasiflat rank of a metric space $X$, we mean the supremal integer $\operatorname{qf.rk}X$ for which there is a quasiisometric embedding $\mathbf{R}^{\operatorname{qf.rk}X}\to X$.
@@
Let $X$ be a coarsely connected coarse median space, and let $\hat{X}$ be an asymptotic cone of $X$; it is automatically a connected median algebra of rank $\operatorname{rk}\hat{X}$. We have $\operatorname{sepdim}\hat{X}=\operatorname{rk}\hat{X}\leq\operatorname{rk}X$. If $X$ is proper and has cocompact isometry group, then the following quantities agree.
@@
  • $\operatorname{qf.rk}X$
  • The supremal $d$ such that there is a quasiisometric embedding $\mathbf{R}^{d}\to X$ that is quasimedian for some coarse median on $X$ realising $\operatorname{rk}X$.
@@
The agreement of $\operatorname{sepdim}\hat{X}$ with $\operatorname{rk}\hat{X}$ is given by Lemmas 2.11 and 2.14. In particular, since separation dimension is a topological property, $\operatorname{rk}\hat{X}$ is independent of the choice of coarse median on $X$. According to [Bow13, Thm 2.3], we have $\operatorname{rk}X\geq\operatorname{rk}\hat{X}$. Also, if $\operatorname{qf.rk}X\geq d$, then $\hat{X}$ contains a bilipschitz copy of $\mathbf{R}^{d}$, and hence $\operatorname{sepdim}\hat{X}\geq\operatorname{qf.rk}X$.
Now suppose that $X$ is proper and has cocompact isometry group. We shall construct a quasiflat of dimension $\operatorname{rk}X$, which will show the equivalence between the first four bulleted items.
Let $d=\operatorname{rk}X$, and let $\mu$ be a coarse median on $X$ with $\operatorname{rk}\mu=d$. Let $(h^{\mu}_{n})$ be a corresponding sequence. For any increasing sequence $(h_{n})$ there is a finite set $A\subset X$ such that there is no median algebra $M$ of rank $d-1$ admitting a $3h_{3^{2d}|A|}$–quasimedian map $\iota:M\to X$ and a map $o:A\to M$ such that $\operatorname{\mathsf{d}}(\iota o(a),a)\leq 3h_{3^{2d}|A|}$ for all $a\in A$. We are free to assume that $h_{n}\geq 3h^{\mu}_{n}$ for all $n$.
Let $N$ be a median algebra of rank $d$ that approximates $A$ with respect to $\mu$. According to [Bow22, Prop. 8.2.4], in any median algebra of rank $d$, the subalgebra generated by a subset is generated by taking medians at most $2d$ times. Hence $|N|\leq 3^{2d}|A|$. Let $\iota^{\mu}$ be the map corresponding to $\mu$, $A$, and $N$. The set $A$ lies in the $h^{\mu}_{|A|}$–neighbourhood of $\iota^{\mu}N$, so the latter, whose cardinality is at most $3^{2d}|A|$, cannot be approximated by any median algebra of rank $d-1$ with error at most $2h_{3^{2d}|A|}$. As $\iota^{\mu}N$ is an $h^{\mu}_{|A|}$–quasisubalgebra, this shows that there is an $h^{\mu}_{|A|}$–quasimedian embedding of a cube $\{0,1\}^{d}\to X$ whose image is $h_{3^{2d}|A|}$–separated. Approximating the images of those cubes with $\mu$ and varying the sequence $(h_{n})$, we conclude that $X$ contains arbitrarily large $h^{\mu}_{2^{d}}$–quasicubes of dimension $d$.
Fix a basepoint $x_{0}\in X$. For each $n$, let $Q_{n}$ be a uniform quasicube in $X$ of dimension $d$ whose vertices are $n$–separated. By an application of [Bow19, Lem. 9.1, Prop. 9.3], for sufficiently large $n$, we can take $Q_{n}$ to be the image of a product of $d$ real intervals of length $n$ under a uniform-quality quasimedian quasiisometric embedding. Write $z_{n}$ for the central point of $Q_{n}$. By cocompactness, we can translate $z_{n}$ into a fixed compact set $C$ containing $x_{0}$ by an isometry $g_{n}$. This gives uniform quasicubes in the sequence of coarse median spaces $(X,g_{n}\mu)$ that are all centred in the compact set $C$.
Now take an (unrescaled) ultralimit. Since $X$ is proper, we have $X=\lim_{\omega}B(x_{0},m)$, where $B(x_{0},m)$ is the ball of radius $m$ centred on $x_{0}$. Choosing $m_{n}$ so that $g_{n}Q_{n}\subset B(x_{0},m_{n})$, we get that $Q=\lim_{\omega}g_{n}Q_{n}\subset X$, because $g_{n}Q_{n}$ is centred in the fixed compact set $C$. Because the $Q_{n}$ are uniform quasiisometric embeddings of increasingly large $d$–cubes, $Q$ is a quasiisometric embedding of $\mathbf{R}^{d}$. We have found the desired $d$–quasiflat in $X$.
@@
The conclusion of Proposition 3.1 can fail without the cocompactness assumption, as shown by the following.
{subhead1}
Let $X\subset(\mathbf{R}^{2},\ell^{1})$ be bounded between the $x$–axis, the line $x=1$, and the graph of the function $x\mapsto\log x$. Since $X$ is affinely convex in the plane, it is a median subalgebra. As a coarse median space, $\operatorname{rk}X=2$, because it contains arbitrarily large squares. However, $\hat{X}$ is a ray.
Let $Y$ be the median metric space constructed from the real line by attaching, for each $n\in\mathbf{N}$, a square of side-length $n$ at the point $n\in\mathbf{R}$, along a vertex. Clearly $\operatorname{qf.rk}Y=1$, but both $Y$ and $\hat{Y}$ contain arbitrarily large squares.
@@ 6.2. Coarse median tubular groups
Let $G$ be a rank-$2$ tubular group. If $G$ admits a coarse median, then $G$ is cocompactly cubulated and virtually compact special.
@@
According to [Bow13, Cor. 8.3], if $G$ admits a coarse median then it has a quadratic isoperimetric function, and so Theorem 6.8 shows that $G$ must be undistorted. Moreover, $G$ has geometric dimension two since it is a graph of two dimensional groups (see [SW79, Prop. 3.6]), so Theorem 3.5 shows that $\operatorname{rk}G\leq 2$. In particular, Theorem 4.3 implies that $G$ cannot have a quasiisometrically embedded 2–RBF. By Lemma 6.9, this means that no vertex group of $G$ can have more than two commensurability classes of incident edge groups. Since $G$ is undistorted, it contains no Baumslag–Solitar subgroups $\operatorname{BS}(m,n)$ with $m\neq\pm n$. Consequently, [Wis14b, Cor 5.10, 5.9] implies that $G$ is cocompactly cubulated and virtually compact special. ∎